# Applying the maximum entropy principle to neural networks enhances multi-species distribution models

### v3

@@ 1 Introduction
In recent years, the rapid growing number of citizen science projects has contributed significantly to the expansion of biodiversity databases, particularly through the collection of presence-only (PO) observations (Callaghan et al., 2022; Bonnet et al., 2020). PO records have been instrumental to improve our understanding of species distributions and help to inform conservation strategies Carvalho et al. (2011); Guisan et al. (2013). However, building reliable Species Distribution Models (SDM) from PO data obtained from opportunistic observations, without a homogeneous sampling protocol, is challenging due to the many sampling biases and errors to handle (Boakes et al. (2010); Bird et al. (2014); Hughes et al. (2021)). Sampling bias correction in SDM becomes even more challenging when SDM become more complex, e.g. based on deep learning (Deneu et al., 2021; Estopinan et al., 2022; Cole et al., 2023). The correction of sampling bias in SDMs based on deep learning has not been addressed to date, yet is an essential step if these methods are to take advantage of PO data.
Maxent (Phillips et al., 2006) is one of the most widely used and effective methods for modelling species distributions based on PO data  (Warren and Seifert, 2011; Elith et al., 2006, 2020; Valavi et al., 2022). Maxent generates a probability distribution across sites representing the sampled area (i.e. background points). The probability distribution is defined as a function of various transformations, hereafter called features, of input environmental variables provided by the user. Maxent’s name arises from the fact that it estimates the probability function by maximising its entropy under constraints on the features. A major issue when calibrating any SDM on PO data is spatial sampling bias, which arises from the clustering of PO records in certain areas, typically of higher accessibility or greater human activity. Such bias can distort SDM outputs, leading to inaccurate species distribution estimates (Yackulic et al., 2013; Fithian et al., 2015; Phillips et al., 2009). To mitigate spatial sampling bias in Maxent, instead of drawing background points uniformly across the area, Phillips et al. (2009) proposed to restrict them to the areas actually sampled, which should reduce the number of false-absences. Specifically, they showed that using occurrences of a group of species, targeted for sampling along with the focal one, to define background points yielded an efficient bias correction (the Target-Group Background correction, hereafter TGB). Furthermore, Maxent is equivalent to a Poisson regression over sites and to inhomogeneous Poisson Point Process (PPP) models  (Renner and Warton, 2013), so that common PPP models of the biased sampling of PO data (Fithian et al., 2015) can be used to show that the TGB correction has theoretical guarantees of robustness to spatial sampling bias under some assumptions (Botella et al., 2020).
@@
Deep learning is a family of data-driven methods that allow fitting arbitrary non-linear functions using neural networks, backpropagation and stochastic gradient descent (Goodfellow et al., 2016; Hornik et al., 1989; LeCun et al., 1989). Unlike other machine learning approaches, deep learning methods allow to simultaneously learn predictive functions while learning an appropriate feature representation, eliminating the need for feature design. Feature design is an important step in traditional SDMs including Maxent (Phillips and Dudík, 2008; Komori et al., 2024), while deep learning-based methods can automatically capture potentially complex and non-linear representations directly from the data (LeCun et al., 2015; Schmidhuber, 2015; Deneu et al., 2021; Estopinan et al., 2022). Learnt features tend to be more predictive and robust the more species are simultaneously used for training (Chen et al., 2017; Botella et al., 2018), leading to a recent interest in deep learning algorithms for multi-species distribution modelling (Kellenberger et al., 2024). Deep learning approaches have been shown to have an edge over other methods in cases where structured and high-dimensional data, such as remote sensing imagery, are used as input variables (Deneu et al., 2021; Estopinan et al., 2024). However, these models are still found to underperform with low-dimensional input variables. After all, they remain susceptible to sampling biases (Zbinden et al., 2024), potentially amplifying them due to their capacity of fitting arbitrary functions, which can lead to inaccurate conclusions and flawed conservation strategies.
In this study, we propose a method that uses the Maxent principle of maximum entropy, with its bias correction capabilities, within a deep learning framework (DeepMaxent) where the features are learnt to predict simultaneously multiple species from PO data. Each species probability distribution across sites is represented as a log-linear function of the joint latent features. Using the equivalence of Maxent and PPPs, we propose a loss function based on the maximum entropy principle of Maxent. We show that computing the loss over any batch of sites preserves the global minimiser of the full loss, so that we can apply the SGD algorithm to train the model and benefit from its regularisation and scalability. We show how the TGB correction can be implemented in DeepMaxent to mitigate spatial sampling bias. We evaluate DeepMaxent on a reference dataset  (Elith et al., 2020) encompassing PO data for training and presence-absence (PA) data for evaluation in six distinct regions and comprising different biological groups. We compare the approach to various state-of-art SDMs, notably Maxent, its TGB correction and other multi-species deep learning based SDM. We also test the robustness of DeepMaxent to various neural network architectures and hyper-parameters for its training.
@@ 2 Materials and Methods
2.1 DeepMaxent: point process for SDM based on neural networks
2.1.1 Point processes and Poisson loss
[…]
PO data, the actual observations, consist of a set of count vectors $y:=\{y_{ij}\}_{i=1,j=1}^{K,N}$ corresponding to the $K$ sites and $N$ species, with $y_{ij}\in\mathbb{N}_{+}$. Most often in PO data, a given $y_{ij}$ is very sparse.
Since the observation process is modelled as an inhomogeneous Poisson point process, a loss function $\mathcal{L}$ is defined to integrate the likelihood of the observed counts as a function of the intensity function $\lambda$ (Cressie, 1993). In this paper, we use the notation $\mathcal{L}_{\mathcal{P}}(\lambda,y)$ to denote the Poisson loss between occurrence counts $y:=\{y_{ij}\}_{i=1,j=1}^{K,N}$ and estimated intensity values $\lambda:=\{\lambda_{ij}\}_{i=1,j=1}^{K,N}$ across the $K$ sites and $N$ species. This can be formulated as:
$\mathcal{L}_{\mathcal{P}}(\lambda,y)=\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(\lambda_{ij}-y_{ij}\log\lambda_{ij}\right).$ (1)
[…]
$\tilde{\lambda}_{ij}=\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}.$ (2)
In this context, $\tilde{\lambda}_{ij}$ can be interpreted as the relative probability of species $j$ being observed at site $i$, given that an occurrence happens somewhere within the $K$ sites. Equivalently for observation counts, the probability function $\tilde{y}_{ij}$ for species $j$, which represents its relative distribution in space, is defined by normalising $y_{ij}$ by the sum of observations $y_{ij}$ over all sites $K$ in the geographical domain $\mathcal{D}$. This normalised intensity $\tilde{y}_{ij}$ for species $j$ at site $i$ is given by :
$\tilde{y}_{ij}=\frac{y_{ij}}{\sum_{k=1}^{K}y_{kj}}.$ (3)
When incorporating this normalisation into the Poisson loss, the original formulation of the Poisson log-likelihood can be reformulated to reflect the normalised intensities. In this context, $\tilde{\lambda}_{ij}$ can be interpreted as the relative probability of species $j$ being observed at site $i$, given that an occurrence happens somewhere within the $K$ sites. Given that $\tilde{\lambda}_{ij}$ now represents a proportion (i.e., a "density-like" measure), the loss function is adjusted to ensure that these normalised values are used consistently. The loss can then be expressed as follows:
$\mathcal{L}_{\mathcal{H}}(\tilde{\lambda},\tilde{y})=-\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(\frac{y_{ij}}{\sum_{k=1}^{K}y_{kj}}\right)\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right).\$ (4)
Note that $\mathcal{L}_{\mathcal{H}}(\tilde{\lambda},\tilde{y})$ represents a cross entropy loss that corresponds to the Maxent loss formulation, which is also equivalent to the conditional negative log-likelihood of a Poisson point process, as previously demonstrated by Renner and Warton (2013). As shown in appendix A.1, it implies that the losses in equations 4 and 1 are equivalent in the space of normalised intensities (i.e. values of $\tilde{\lambda}$). However, the Poisson loss introduces a constant factor in the intensity that can affect the learning of the parameters determining $\tilde{\lambda}$. In the multi-species case, however, the per-species normalisation of the PO labels $y_{ij}$ according to the total number of observations $\sum_{k=1}^{K}y_{kj}$ leads to all species having the same weight in the total loss. This can lead to learning a joint latent representation (see below) that is overwhelmed by the noisy learning signal from rare species. To address this problem, we propose to weight the loss per species proportionally to their number of PO occurrences $w_{j}:=\sum_{i=1}^{K}y_{ij}$, which leads to:
$\mathcal{L}_{\mathcal{H,W}}(\tilde{\lambda},y)=-\frac{1}{KN}\sum_{i=1}^{K}w_{j}\sum_{j=1}^{N}\left(\frac{y_{ij}}{\sum_{k=1}^{K}y_{kj}}\right)\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)=-\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}y_{ij}\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)$ (5)
Note that if the intensity of each species was based on pre-determined covariates, this weighting wouldn’t have any impact on the estimated intensities. However, given that we learn a joint features representation between species (as introduced below), the model can calibrate this representation to favour performance on certain species at the expense of others during learning.
2.1.2 Feature extraction based on neural network
@@
In Maxent, the intensity value in a given site is a function of a vector of environmental variables $x\in\mathbb{R}^{P}$ in that site. More precisely, this intensity is defined as a log-linear function of a feature vector $f(x)$, composed of pre-determined transformations of $x$. We extend the principle of Maxent by replacing its feature vector $f(x)$ with a feature extractor instantiated as a neural network $g:\mathbb{R}^{P}\mapsto\mathbb{R}^{C}$ parametrised by $\theta$, where $C$ is the dimensionality of the last hidden layer, the joint latent representation that is common to all species, as illustrated in Figure 1. The intensity function $\lambda_{j}(x)$ of species $j$ in DeepMaxent is then given by:
$\lambda_{j}(x)=\exp\big{(}\sum_{c=1}^{C}\gamma_{jc}g(x;\theta)_{c}+b_{j}\big{)},$ (6)
where $\gamma\in\mathbb{R}^{N\times C}$ and $b\in\mathbb{R}^{N}$ can be treated the weights and the bias term of an additional linear layer that maps from $\mathbb{R}^{C}$ to $\mathbb{R}^{N}$, corresponding to the number of species. The function $g$ can automatically learn complex, non-linear relationships between environmental variables and the presence of multiple species from the data, potentially enabling the model to identify environmental patterns not considered by traditional approaches. A key advantage of DeepMaxent is its scalability to many species. DeepMaxent computes a single feature vector $g(x)$ per site for all species, so adding more species does not increase the computational cost of computing $g(x)$. In addition, all the species contribute to learning $g$, which means that species with very few observations can benefit from the learning signal provided by other species. These properties are interesting when $g$ is designed to have more capacity, e.g. when it includes several hidden layers or if the neural network takes high dimensional input variables such as spatio-temporal data.
Figure 1: The residual neural network to estimate the intensity $\lambda$ from variable input ($x_{0}$), where $P$ as is the variable number input, where $C$ is the number of hidden layer nodes, and $N$ denotes the number of species (or target categories). The illustrated case involves two hidden layers. In the special case where there is only one hidden layer, no residual addition is applied.
DeepMaxent is architecture-agnostic, and in this work we chose a $g$ based on a residual neural network architecture (see Figure 1), which takes the vector of environmental variables $x$ as input. $x$ is transformed by a sequence of hidden layers $g=g_{1}\circ g_{2}\circ\dots\circ g_{L}$, whose number $L$ and size are the main hyper-parameters. Residual connections between adjacent hidden layers were incorporated introducing shortcuts that allow the input of a layer to bypass the non-linear transformations and be directly added to the output ($x_{l+1}=x_{l}+g_{l}(x_{l})$). The output of the last hidden layer ($g(x)$), used as our latent feature vector, leads to the output layer, which is composed of each species intensity measure $\lambda_{j}(x)$.

### v4

@@ 1 Introduction
In recent years, the rapid growth of citizen science projects has contributed significantly to the expansion of biodiversity databases. Among the different types of data collected, a large amount consists of presence-only (PO) observations (Callaghan et al., 2022; Bonnet et al., 2020). PO records have been instrumental in improving our understanding of species distributions and helping inform conservation strategies (Carvalho et al., 2011; Guisan et al., 2013).
Maxent (Phillips et al., 2006) is one of the most widely used and effective methods for Species Distribution Model (SDM) based on PO data  (Warren and Seifert, 2011; Elith et al., 2006, 2020; Valavi et al., 2022). Maxent generates a relative probability of species occurrence across space as a function of environmental variables. This function is applied to various predefined transformations of input environmental variables. Maxent’s output can be interpreted as the probability of observing the species in each site relative to the other sites and knowing that it has been observed once. This probability estimate of Maxent is actually equivalent to the one that is derived from a related Poisson regression or a spatially discretized Poisson process (Renner and Warton, 2013). Maxent’s name arises from the fact that its formulation leads to finding, among all solutions that fit the PO training data, the one that maximizes the entropy of this spatial probability distribution. This behaviour, in which spatially smooth solutions are favoured for regions with sparse PO data, participated to Maxent’s robust performances against many other SDM methods under such data regimes and in diverse contexts (Elith et al., 2006; Valavi et al., 2022).
Yet, a major issue when calibrating SDMs using PO data is spatial sampling bias, which leads to clustering of PO records in areas with high sampling effort, typically of higher accessibility or greater human activity. Such bias can distort SDM outputs, leading to inaccurate species distribution estimates (Yackulic et al., 2013; Fithian et al., 2015; Phillips et al., 2009).
A wide range of strategies have been proposed to correct for spatial sampling bias in SDMs, including methods based on background points manipulation, spatial filtering of records or explicit bias modelling (Phillips et al., 2009; Boria et al., 2014; Fithian et al., 2015). Specifically for Maxent, (Phillips et al., 2009) proposed the Target-Group Background correction (hereafter TGB), which restricts the background sites used by Maxent to those where at least one species was observed among a Target Group of species. This Target Group should contain species being sampled along with the focal one. Phillips et al. (2009) evaluated the TGB correction with Maxent on a large standardized dataset. This correction was also robust in later studies (Fourcade et al., 2014). Besides its simplicity and empirical robustness, the TGB correction also comes with theoretical guarantees. Indeed, with the assumptions that all species occurrences are drawn from independent Poisson point processes thinned by a same sampling bias, as it is often assumed (Fithian et al., 2015; Botella et al., 2021), and that the TG species cumulated intensities are constant, TGB yields an unbiased estimate of the species relative intensity across sites (Botella et al., 2020), which then applies within Maxent (Renner and Warton, 2013). More recent studies have proposed other bias corrections in the context of deep learning based SDMs, by adapting the loss function to weight presences and background points (Zbinden et al., 2024; Gillespie et al., 2024).
@@
Feature design, i.e. the definition of pre-defined transformations (features) of the input variables, is an important step in traditional SDMs, including Maxent (Phillips and Dudík, 2008; Komori et al., 2024). Deep learning is a family of data-driven methods that removes the need for feature design by allowing the model to learn arbitrary non-linear features from the data using neural networks, backpropagation and stochastic gradient descent (Goodfellow et al., 2016; Hornik et al., 1989; LeCun et al., 1989). While other approaches are also capable of learning non-linear relationships, such as Generalized Additive Models (GAM), Multivariate Adaptive Regression Splines (MARS) or Boosted Regression Trees (BRT), and have been used for SDMs (Phillips et al., 2009), deep learning methods offer a broader and more flexible class of models that can automatically learn rich, hierarchical features while integrating efficient regularisation strategies to mitigate overfitting  (LeCun et al., 2015; Schmidhuber, 2015). Additionally, like other multi-species SDMs do, although typically restricted to the linear case (Ovaskainen et al., 2017b; van der Veen et al., 2023). Deep learning based SDM (deepSDMs) can also learn shared features to simultaneously predict multiple species distributions. Furthermore, these learnt features tend to be more predictive and robust the more species are included  (Chen et al., 2017; Botella et al., 2018), leading to a recent interest in deep learning for multi-SDMs  (Kellenberger et al., 2024). In addition, deepSDM architectures can capture predictive features from structured and high-dimensional input data, such as remote sensing imagery  (Deneu et al., 2021; Estopinan et al., 2024; Deneu et al., 2021; Estopinan et al., 2022). In spite of this, deepSDM performance have remained limited so far when using low-dimensional input variables  (Zbinden et al., 2024). Besides, they remain susceptible to sampling biases  (Zbinden et al., 2024), and potentially amplifying them due to their capacity of fitting arbitrary functions.
In this study, we propose DeepMaxent, a method that combines the Maxent principle of maximum entropy with the data-driven feature extraction capabilities of deep learning methods. The DeepMaxent model uses PO data to jointly learn shared latent features and the functions that map from these features to the probability distribution across sites for each species. We propose a loss function, hereafter the DeepMaxent loss, that generalizes Maxent for modelling the probability function with a wider class of functions, including neural networks, preserving the equivalence with the Poisson regression loss (Renner and Warton, 2013). In contrast to loss functions often used for deepSDMs, which attempt at modelling the relative probability of each species given a site, DeepMaxent aims at modelling the probability of selecting each site for an observation given a species, as done in Maxent (see Figure 1). Note that the latter is an easier objective, since modelling the relative probability of each species given a site requires capturing the relative abundances between species. Unlike the original Maxent loss, which is optimised accounting for the whole dataset at every optimisation step, we adopt a mini batch-based approach to ensure scalability in terms of data and model size. We show that the global minimizer of such loss is the same as the one using the full dataset to inform each optimisation step. Similarly to Maxent, we show how the TGB correction can be implicitly incorporated into DeepMaxent to efficiently mitigate spatial sampling bias. We evaluate DeepMaxent and compare it to alternative methods on two reference benchmarks (Elith et al., 2020; Picek et al., 2025), both encompassing PO data for SDM training and presence-absence (PA) data for evaluation. We carried an extensive comparative evaluation on the NCEAS dataset  (Elith et al., 2020), comprising six distinct regions and different biological groups. We compare DeepMaxent to alternative loss functions (Poisson, Cross-Entropy across species, Binary Cross Entropy) with or without the TGB correction, and to various state-of-the-art SDMs, notably Maxent and other multi-species deep learning based SDM. We also conduct sensitivity and ablation studies on this dataset to assess the importance of DeepMaxent’s hyper-parameters and components. We conduct additional experiments on the GeoPlant dataset (Picek et al., 2025), including an independent comparison of DeepMaxent and Maxent, and an illustration of how DeepMaxent can leverage remote sensing input data with an assessment of the related performance gains compared to tabular climatic data.
@@ 2 Materials and Methods
Figure 1: Illustration of three species distribution modelling approaches (Maxent, DeepMaxent, and a common cross-entropy). Maxent relies on handcrafted environmental features and trains an independent model for each species. DeepMaxent uses a single deep learning model to predict multiple species simultaneously, incorporating batch normalisation across samples to standardize feature representations. In contrast, the commonly used cross-entropy loss approach applies normalisation along the species dimension, focusing on predicting the presence of species at each site rather than modelling species jointly.
@@
2.1 DeepMaxent: Maximum entropy principle for SDMs based on neural networks
2.1.1 A generalization of Maxent’s loss function
We introduce here the statistical model underlying the investigated methods. We consider a geographic domain $\mathcal{D}\subset\mathbb{R}^{2}$ composed of $K$ non-overlapping spatial areas $d_{1},...,d_{K}\subset\mathcal{D}$ (e.g. a regular mesh), hereafter called sites. We consider $N\in\mathbb{N}^{*}$ species indexed by $j$, and note $y_{ij}\in\mathbb{N}$ the count of PO observations for species $j$ in site $i\in[1,K]$. For each site $i\in[1,K]$, we have covariates $\mathbf{x}_{i}\in\mathbb{R}^{P}$ typically encoding environmental factors (e.g. climate or soil properties), as well as positive intensity values $\lambda_{ij}>0$ for and species $j$. We further assume $\lambda_{ij}$ to be a parametric function of the covariates of the form $\lambda_{ij}=\lambda_{j}(\mathbf{x}_{i})$, where we explicitly define $\lambda_{j}:\mathbb{R}^{P}\to\mathbb{R}_{+}$ as a positive intensity function mapping site-level covariates to expected counts. This function is then parametrized as $\lambda_{j}(\mathbf{x}_{i})=\exp(b_{j}+f_{\theta_{j}}(\mathbf{x}_{i}))$ with real parameters $b_{j},\theta_{j}$. The function $f_{\theta_{j}}$ can be an arbitrary mapping $f_{\theta_{j}}:\mathbb{R}^{P}\to\mathbb{R}$. In a Poisson regression setting, the counts $y_{ij}$ are modelled as Poisson random variables $y_{ij}\sim\mathcal{P}(\lambda_{ij})$ where these counts are independent between sites and species. This Poisson regression model notably corresponds to the discrete approximation of a spatial Inhomogeneous Poisson Process (IPP), and is very often used for SDMs (Renner et al., 2015) with a known equivalence to Maxent when $f_{\theta_{j}}$ is linear in $x_{i}$ (Renner and Warton, 2013). Such discretization is common and handful to avoid the computational burden of fitting an IPP on a continuous spatial domain where covariates vary at high resolution, and it makes particular sense when covariates come from a stack of geographic rasters.
We note for convenience the intensity values (i.e. model predicted counts) across all sites and species by the matrix $\boldsymbol{\Lambda}=\{\lambda_{ij}\}_{i=1,j=1}^{K,N}\in\mathbb{R}_{+}^{K\times N}$, and similarly $\mathbf{Y}=\{y_{ij}\}_{i=1,j=1}^{K,N}\in\mathbb{N}_{+}^{K\times N}$ for the PO count data. The negative log-likelihood of the above described Poisson regression, hereafter the Poisson loss, is written in equation 1.
$\mathcal{L}_{\mathcal{P}}(\boldsymbol{\Lambda},\mathbf{Y})=\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(\lambda_{ij}-y_{ij}\log\lambda_{ij}\right),$ (1)
The parameters’ estimates for the Poisson loss (i.e. the maximum likelihood estimates of the Poisson regression) can then be noted $(\hat{b}_{1}^{\mathcal{P}},...,\hat{b}_{N}^{\mathcal{P}},\hat{\theta}_{1}^{\mathcal{P}},...,\hat{\theta}_{N}^{\mathcal{P}})=\underset{b_{1},...,b_{N},\theta_{1},...,\theta_{N}}{\text{argmin}}\mathcal{L}_{\mathcal{P}}(\boldsymbol{\Lambda},\mathbf{Y})$. It is noteworthy that, due to the equivalence from Renner and Warton (2013), the sites included in the terms of equation 1 can be interpreted as the background sites (or points) of Maxent (Phillips et al., 2009), as developed in the dedicated section below.
We introduce in equation 2 a new loss $\mathcal{L}_{\mathcal{H}}$ derived from the Poisson loss of equation 1, named DeepMaxent loss. We can see from this equation that the DeepMaxent loss is a modification of the Poisson loss where the counts $y_{ij}$ and intensity values $\lambda_{ij}$ are normalised by their sum over sites, and each species term is weighted by the species total PO count. The loss measures the discrepancy between the observed and predicted probability distributions of PO across sites for each species. Indeed, equation 2 corresponds to a weighted sum of species-wise cross-entropy losses (or Kullback–Leibler divergences) between the empirical and predicted distribution of PO records over sites.
$\mathcal{L}_{\mathcal{H}}(\boldsymbol{\Lambda},\mathbf{Y})=-\frac{1}{KN}\sum_{j=1}^{N}\sum_{i=1}^{K}y_{ij}\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)=-\frac{1}{KN}\sum_{j=1}^{N}(\sum_{k=1}^{K}y_{kj})\sum_{i=1}^{K}\left(\frac{y_{ij}}{\sum_{k=1}^{K}y_{kj}}\right)\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)\\$ (2)
We show in Appendix A.1 that the DeepMaxent loss generalizes the loss of Maxent to intensity models $\lambda_{ij}=\exp(b_{j}+f_{\theta_{j}}(x_{i}))$ which are not necessarily log-linear, e.g. when $f$ is built on neural networks, and obviously to multiple species. That is, if $\lambda_{ij}$ is log-linear, the DeepMaxent loss is equivalent to the one of Maxent. Furthermore, Appendix A.1 shows that we preserve the equivalence with the Poisson estimate Renner and Warton (2013) for non log-linear intensities: The global minimizer of the DeepMaxent loss for the parameters of the probability distribution across sites $(\hat{\theta}^{\mathcal{H}}_{1},...,\hat{\theta}^{\mathcal{H}}_{N})$ is equal to the one of the Poisson loss $(\hat{\theta}^{\mathcal{P}}_{1},...,\hat{\theta}^{\mathcal{P}}_{N})$ of equation 1, and the difference with the Poisson loss is that the latter provides an estimate for $b_{j}$. Although $b_{j}$ is not identifiable under the DeepMaxent loss, as it cancels out in equation 2, we consider it for consistency across all losses introduced and tested below, including the Poisson loss.
2.1.2 Feature extraction using neural networks
@@
In Maxent, the intensity is defined as a log-linear function of a feature vector $f(\mathbf{x})$, composed of pre-determined transformations of $\mathbf{x}$. We extend the principle of Maxent by replacing $f(\mathbf{x})$ with a feature extractor instantiated as a neural network $g_{\theta}:\mathbb{R}^{P}\mapsto\mathbb{R}^{C}$ parametrised by $\theta$, where $C$ is the dimensionality of the last hidden representation. The output $g_{\theta}(\mathbf{x})\in\mathbb{R}^{C}$ is a shared latent representation across all species, as illustrated in Figure 5. The intensity function $\lambda_{j}(\mathbf{x})$ of species $j$ in DeepMaxent is then given by:
$\lambda_{j}(\mathbf{x})=\exp\big(\sum_{c=1}^{C}\gamma_{jc}g_{\theta}(\mathbf{x})_{c}+b_{j}\big),$ (3)
where $\gamma_{j}\in\mathbb{R}^{C}$ is the species-specific weight vector and $b_{j}\in\mathbb{R}$ is the species-specific intercept. Collectively, the weight vectors for all species form the matrix $\Gamma\in\mathbb{R}^{N\times C}$, and the intercepts form the vector $\mathbf{b}\in\mathbb{R}^{N}$. The function $g_{\theta}$ can automatically learn complex, non-linear relationships between environmental variables and the presence of multiple species from the data, potentially enabling the model to identify environmental patterns that cannot be captured by a linear mapping. With this multi-species architecture, similarly to many earlier deep learning based SDM implementations, we have one shared $g_{\theta}$ across all species, which offers a key computational advantage compared to training one DeepMaxent model per species, given the complexity of $g_{\theta}$. Indeed, at each gradient descent step, $g_{\theta}$ and its gradient are computed only once per site, independent of the number of species. Besides, sharing this representation of the environment across species yielded more robust performances in deep learning based SDMs (Botella et al., 2018), especially for data-poor species.